# Evaluating LLM-corrupted Crowdsourcing Data Without Ground Truth

### v1

@@ Appendix A Preliminary
What do we know about the problem? Under the assumption that there are no cheap signals, the peer prediction literature suggests that estimating the mutual information between $X_{i}$ and $X_{j}$ is an information monotone metric [Kong and Schoenebeck, 2019a]. There are various ways to estimate the mutual information using agents’ responses to multiple questions [Kong, 2020a, Schoenebeck and Yu, 2020] and many information monotone metrics that can be explained using mutual information [Dasgupta and Ghosh, 2013, Shnayder et al., 2016]. In this section, we discuss a simple yet intuitive mechanism called the correlated agreement (CA) mechanism [Shnayder et al., 2016].
@@
The CA mechanism rewards agents when they “agree” on the same question and penalizes them when they “agree” on distinct questions. Instead of directly checking agreement, the mechanism redefines agreement based on the correlation between signals. A pair of signals $\sigma,\sigma^{\prime}\in\Sigma$ is correlated agreed if $\Pr(X_{1}=\sigma,X_{2}=\sigma^{\prime})-\Pr(X_{1}=\sigma)\Pr(X_{2}=\sigma^{\prime})>0$, meaning that $\sigma$ and $\sigma^{\prime}$ are more likely to appear on the same question rather than on distinct questions. In particular, the mechanism utilizes the delta matrix $\Delta_{\sigma,\sigma^{\prime}}=\Pr(X_{1}=\sigma,X_{2}=\sigma^{\prime})-\Pr(X_{1}=\sigma)\Pr(X_{2}=\sigma^{\prime})$ to design the scoring function $T_{\Delta}=\text{Sign}(\Delta)$ where $\text{Sign}(x)=1$ if $x>0$ and 0 otherwise.
@@
In practice, the mechanism can be implemented in two ways depending on the information available to the principal. First, if $T_{\Delta}$ is known, the mechanism can be directly implemented, as shown in Algorithm 2. This case happens if the information structure is simple, e.g. when questions have binary answers, $T_{\Delta}$ is likely to be a $2\times 2$ identical matrix. Second, if $T_{\Delta}$ is unknown, the mechanism can first estimate the joint distribution, denoted as $\widetilde{\Pr}(X_{1},X_{2})$ using the report matrix $\hat{X}$. Then, an estimate of $T_{\Delta}$ can be computed using the estimated Delta matrix $\tilde{\Delta}_{\sigma,\sigma^{\prime}}=\widetilde{\Pr}(X_{1}=\sigma,X_{2}=\sigma^{\prime})-\widetilde{\Pr}(X_{1}=\sigma)\Pr(X_{2}=\sigma^{\prime})$. Next, Algorithm 2 can be applied.
@@
Algorithm 2 The Correlated Agreement Mechanism [Shnayder et al., 2016]
Appendix B Proofs
B.1 Sufficiency of the Assumptions
@@
$\geq$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\left(\Pr(z_{i}\mid z)\Pr(x_{i}\mid z)-\epsilon_{i}\sum_{y\in\Sigma}\Pr(x_{i}\mid y,z)\right)\;\Pr(x_{j},z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z).$ (By Lemma B.1)
@@
We focus on the second term. First note that $T_{\Delta}\leq 1$. Therefore, the second term is lower bounded by $-\epsilon_{i}\sum_{y,z_{i}\in\Sigma}\sum_{x_{j},z_{j}\in\Sigma}\Pr(x_{j},z_{j}\mid z)\sum_{x_{i}\in\Sigma}\Pr(x_{i}\mid y,z)$. Marginalizing over $x_{j},z_{j}$ and $x_{i}$ and observing that $y,z_{i}\in\Sigma$, the quality is lower-bounded by $\epsilon_{i}|\Sigma|^{2}$. This means the original quantity is lower-bounded by
@@
$\qquad\qquad\cdot T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)-\epsilon_{i}|\Sigma|^{2}$ (By Lemma B.1)
@@
$\leq$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\sum_{y\in\Sigma}\Pr(y\mid z)\Pr(x_{i},x_{j}\mid y,z)\left(\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)+\epsilon+\frac{\epsilon_{i}+\epsilon_{j}}{\Pr(y|z)}\right)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z).$ (By Lemma B.2)
@@
Eq. 3 $=\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)\sum_{y\in\Sigma}\Pr(x_{i},x_{j},y\mid z)$
@@
Eq. 4 $=\epsilon\;\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)\;\sum_{y\in\Sigma}\Pr(x_{i},x_{j},y\mid z)$
@@
Eq. 5 $=(\epsilon_{i}+\epsilon_{j})\;\sum_{y\in\Sigma}\sum_{z_{i},z_{j}\in\Sigma}\sum_{x_{i},x_{j}\in\Sigma}\Pr(x_{i},x_{j}\mid y,z)T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$
@@
Finally, combining Eq. 2 and Eq. 6,
@@
Lastly, we show that if agent $i$ plays a naive strategy $\nu_{i}$, the expected score is strictly smaller than truth-telling. Note that because a naive strategy is a convex combination between truth-telling $\tau$ and an uninformative strategy $\mu_{i}$, it suffices to show that the expected score of any uninformative strategy is strictly smaller than truth-telling. Because $\mu_{i}(X_{i},Z_{i})$ is independent of $X_{i}$, we can reoder the summations in Eq. 7 in the following way. Let $\mu=\mu_{i}(X_{i},Z_{i})$.
@@
$\sum_{x_{i}\in\Sigma}\left(\Pr(x_{i},x_{j}\mid z)-\Pr(x_{i}\mid z)\Pr(x_{j}\mid z)\right)+\hat{\epsilon}$ (by Eq. 7)
@@
Note that the expected score under truth-telling is exactly $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$. Therefore, whenever $I_{\text{TVD}}(X_{i};X_{j}\mid Z)>\hat{\epsilon}$, we have $S_{i}(\tau,\tau)>S_{i}(\mu_{i},\theta_{j})$ for any uninformative strategy $\mu_{i}$ and any strategy $\theta_{j}$. This completes the proof. ∎
Proof of Lemma B.1.
$\Pr(X_{i},Z_{i}\mid Z)$ $=\sum_{y\in\Sigma}\Pr(Y=y\mid Z)\Pr(X_{i},Z_{i}\mid Y=y,Z)$ (By the law to total probability)
@@
Proof of Lemma B.2.
$\left|\Pr(Z_{1},Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Z)\Pr(Z_{2}\mid Z)\right|$
$=$ $\left|\Pr(Z_{1},Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)+\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Z)\Pr(Z_{2}\mid Z)\right|$
$\leq$ $\left|\Pr(Z_{1},Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)\right|+\left|\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Z)\Pr(Z_{2}\mid Z)\right|$ (Triangle inequality)
$\leq$ $\epsilon+\left|\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Z)\Pr(Z_{2}\mid Z)\right|.$ (4.2)
@@
$\left|\Pr(Z_{1}\mid Y,Z)\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{1}\mid Z)\Pr(Z_{2}\mid Z)\right|$
$=$ $\left|\Pr(Z_{1}\mid Y,Z)(\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{2}\mid Z))+\Pr(Z_{2}\mid Z)(\Pr(Z_{1}\mid Y,Z)-\Pr(Z_{1}\mid Z))\right|$
$\leq$ $\left|\Pr(Z_{1}\mid Y,Z)(\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{2}\mid Z))\right|+\left|\Pr(Z_{2}\mid Z)(\Pr(Z_{1}\mid Y,Z)-\Pr(Z_{1}\mid Z))\right|$ (Triangle inequality)
$\leq$ $\left|\Pr(Z_{2}\mid Y,Z)-\Pr(Z_{2}\mid Z)\right|+\left|\Pr(Z_{1}\mid Y,Z)-\Pr(Z_{1}\mid Z)\right|$
$=$ $\frac{1}{\Pr(Y\mid Z)}\;\left|\Pr(Z_{2},Y\mid Z)-\Pr(Y\mid Z)\;\Pr(Z_{2}\mid Z)\right|+\frac{1}{\Pr(Y\mid Z)}\left|\Pr(Z_{1},Y\mid Z)-\Pr(Y\mid Z)\;\Pr(Z_{1}\mid Z)\right|$
@@ B.2 Necessity of the Assumptions
We show that this score is strictly smaller than the expected score when agents report 00 if and only if $X_{i}=Z_{i}=0$ (and $X_{j}=Z_{j}=0$) and report $1$ otherwise. Let $\theta$ be the strategy such that $\theta(k,l|z)=0$ if and only if $k=l=0$ for any $z$. We can write down the joint distribution and the product of marginal distributions for $\hat{X}_{i}$ and $\hat{X}_{j}$.
$\Pr(\hat{X}_{i},\hat{X}_{j}\mid y,z)=\begin{pmatrix}2d&0\\ 0&1-2d\end{pmatrix}\qquad\Pr(\hat{X}_{i},\hat{X}_{j}\mid y,z)=\begin{pmatrix}4d^{2}&2d(1-2d)\\ 2d(1-2d)&(1-2d)^{2}\end{pmatrix}\quad\text{for any $y,z\in\{0,1\}$.}$
@@
$=S_{i}(\tau,\tau)+8d^{2}.$
B.3 Information Monotonicity under Naive Strategies
@@
A naive strategy is a convex combination of truth-telling and an uninformative strategy. Therefore, it suffices to show that the expected score of truth-telling is strictly higher than that of any uninformative strategy.
An uninformative strategy is essentially a mixture of two types of strategy. First, $\mu_{i}(X_{i},Z_{i})$ is independent of both $X_{i}$ and $Z_{i}$. In this case, we have shown in the proof of Theorem 4.3 that the expected score is 0, strictly less than the expected score of truth-telling, which is $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$.
Second, $\mu_{i}(X_{i},Z_{i})$ depends only on $Z_{i}$; let $\mu_{i}(X_{i},Z_{i})=\hat{Z}_{i}$. Conditioned on $Z$, the Markov chain $Z_{i}\rightarrow\hat{Z}_{i}$ holds. Now, fix agent $j$’s naive strategy and let her report be $\hat{X}_{j}$. By the data processing inequality [Shannon, 1948], we have:
@@
Next, suppose agent $i$ reporting $Z_{i}$, which is the score-maximizing naive strategy regardless of agent $j$’s strategy. For any naive strategy such that $\hat{X}_{j}=\nu_{j}(X_{j},Z_{j})$ that agent $j$ plays, applying the data processing inequality to agent $j$’s report yields:
@@
Therefore, 4.5 is sufficient for information monotonicity under naive strategies. ∎
B.4 Learning the Scoring Function
Proof of 4.7.

### v2

@@ Appendix B Preliminary
Under the assumption that there are no cheap signals, the peer prediction literature suggests that estimating the mutual information between $X_{i}$ and $X_{j}$ is an information monotone metric FrameworkKong2019. There are various ways to estimate the mutual information using agents’ responses to multiple questions dmiKong2020, schoenebeck2020learning and many information monotone metrics that can be explained using mutual information dasgupta2013crowdsourced, shnayder2016informed. In this section, we discuss a simple yet intuitive mechanism called the correlated agreement (CA) mechanism [shnayder2016informed].
@@
The CA mechanism rewards agents when they “agree” on the same question and penalizes them when they “agree” on distinct questions. Instead of directly checking agreement, the mechanism redefines agreement based on the correlation between signals. A pair of signals $\sigma,\sigma^{\prime}\in\Sigma$ is correlated agreed if $\Pr(X_{i}=\sigma,X_{j}=\sigma^{\prime})-\Pr(X_{i}=\sigma)\Pr(X_{j}=\sigma^{\prime})>0$, meaning that $\sigma$ and $\sigma^{\prime}$ are more likely to appear on the same question rather than on distinct questions. In particular, the mechanism utilizes the delta matrix $\Delta_{\sigma,\sigma^{\prime}}=\Pr(X_{i}=\sigma,X_{j}=\sigma^{\prime})-\Pr(X_{i}=\sigma)\Pr(X_{j}=\sigma^{\prime})$ to design the scoring function $T_{\Delta}=\operatorname{sign}(\Delta)$ where $\operatorname{sign}(x)=1$ if $x>0$ and 0 otherwise.
@@
In practice, the mechanism can be implemented in two ways depending on the information available to the principal. First, if $T_{\Delta}$ is known, the mechanism can be directly implemented, as shown in Algorithm 1. This case happens if the information structure is simple, e.g. when questions have binary answers, $T_{\Delta}$ is likely to be a $2\times 2$ identical matrix. Second, if $T_{\Delta}$ is unknown, the mechanism can first estimate the joint distribution, denoted as $\widetilde{\Pr}(\hat{X}_{i},\hat{X}_{j})$ using the report matrix $\hat{X}$. Then, an estimate of $T_{\Delta}$ can be computed using the estimated Delta matrix $\tilde{\Delta}_{\sigma,\sigma^{\prime}}=\widetilde{\Pr}(\hat{X}_{i}=\sigma,\hat{X}_{j}=\sigma^{\prime})-\widetilde{\Pr}(\hat{X}_{i}=\sigma)\widetilde{\Pr}(\hat{X}_{j}=\sigma^{\prime})$. Next, Algorithm 1 can be applied.
@@
Algorithm 1 The Correlated Agreement Mechanism [shnayder2016informed]
Appendix C Details of the Conditioned Correlated Agreement Mechanism
After learning the scoring function, the conditioned CA mechanism scores agents according to Algorithm 2. At a high level, the mechanism will iteratively fix a signal $k\in\Sigma$ and look at the questions on which $Z_{q}=k$. Let $M_{Z}(k)$ be the set of questions on which $Z=k$. Then, an agent $i$ will receive a higher score if she agrees with another agent $j$ on the same question more often than on two distinct questions conditioning that the questions are drawn from $M_{Z}(k)$. This measures the additional correlation between two independent agents conditioned on $Z$.
for agent $i\in[n]$ do
   Let $M_{i}$ be the set of questions answered by agent $i$;
   Initialize $s_{i}=0$;
 for $k\in\Sigma$ do
      If $M_{i}\cap M_{Z}(k)=\emptyset$, continue;
      Initialize $x=0$;
    for question $q\in M_{i}\cap M_{Z}(k)$ do
         Randomly select a peer $j$ who answers $q$;
       $x\mathrel{+}=T_{\tilde{\Delta}}(\hat{X}_{i,q},\hat{X}_{j,q},k)$;
         Randomly select a question $q^{\prime}\in M_{j}\cap M_{Z}(k)$;
       $x\mathrel{-}=T_{\tilde{\Delta}}(\hat{X}_{i,q},\hat{X}_{j,q^{\prime}},k)$
    $s_{i}\mathrel{+}=x\cdot\tilde{P}(Z=k)/|M_{i}\cap M_{Z}(k)|$
return $s=(s_{1},\ldots,s_{n})$
Algorithm 2 The conditioned CA mechanism
Appendix D Proofs
D.1 Sufficiency of the Assumptions
@@
$\geq$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\left(\Pr(z_{i}\mid z)\Pr(x_{i}\mid z)-\epsilon_{i}\sum_{y\in\Sigma}\Pr(x_{i}\mid y,z)\right)\;\Pr(x_{j},z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z).$ (By Lemma D.1)
@@
We focus on the second term. First note that $T_{\Delta}\leq 1$. Therefore, the second term is lower bounded by $\epsilon_{i}\sum_{y,z_{i}\in\Sigma}\sum_{x_{j},z_{j}\in\Sigma}\Pr(x_{j},z_{j}\mid z)\sum_{x_{i}\in\Sigma}\Pr(x_{i}\mid y,z)$. Marginalizing over $x_{j},z_{j}$ and $x_{i}$ and observing that $y,z_{i}\in\Sigma$, the quality is lower-bounded by $\epsilon_{i}|\Sigma|^{2}$. This means the original quantity is lower-bounded by
@@
$\qquad\qquad\cdot T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)-\epsilon_{i}|\Sigma|^{2}$ (By Lemma D.1)
@@
$\leq$ $\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\sum_{y\in\Sigma}\Pr(y\mid z)\Pr(x_{i},x_{j}\mid y,z)\left(\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)+\epsilon+\frac{\epsilon_{i}+\epsilon_{j}}{\Pr(y|z)}\right)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z).$ (By Lemma D.2)
@@
Equation 3 $=\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}\Pr(z_{i}\mid z)\Pr(z_{j}\mid z)\;T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)\sum_{y\in\Sigma}\Pr(x_{i},x_{j},y\mid z)$
@@
Equation 4 $=\epsilon\;\sum_{x_{i},x_{j},z_{i},z_{j}\in\Sigma}T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)\;\sum_{y\in\Sigma}\Pr(x_{i},x_{j},y\mid z)$
@@
Equation 5 $=(\epsilon_{i}+\epsilon_{j})\;\sum_{y\in\Sigma}\sum_{z_{i},z_{j}\in\Sigma}\sum_{x_{i},x_{j}\in\Sigma}\Pr(x_{i},x_{j}\mid y,z)T_{\Delta}(\theta_{i}(x_{i},z_{i}),\theta_{j}(x_{j},z_{j}),z)$
@@
Finally, combining Equation 2 and Equation 6,
@@
Lastly, we show that if agent $i$ plays a lazy-reporting strategy $\nu_{i}$, the expected score is strictly smaller than truth-telling. Note that because a lazy-reporting strategy is a convex combination between truth-telling $\tau$ and a no-effort strategy $\mu_{i}$, it suffices to show that the expected score of any no-effort strategy is strictly smaller than truth-telling. Because $\mu_{i}(X_{i},Z_{i})$ is independent of $X_{i}$, we can reoder the summations in Equation 7 in the following way. Let $\mu=\mu_{i}(X_{i},Z_{i})$.
@@
$\sum_{x_{i}\in\Sigma}\left(\Pr(x_{i},x_{j}\mid z)-\Pr(x_{i}\mid z)\Pr(x_{j}\mid z)\right)+\hat{\epsilon}$ (by Equation 7)
@@
Note that the expected score under truth-telling is exactly $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$. Therefore, whenever $I_{\text{TVD}}(X_{i};X_{j}\mid Z)>\hat{\epsilon}$, we have $S_{i}(\tau,\tau)>S_{i}(\mu_{i},\theta_{j})$ for any no-effort strategy $\mu_{i}$ and any strategy $\theta_{j}$. This completes the proof. ∎
Proof of Lemma D.1.
$\Pr(X_{i},Z_{i}\mid Z)$ $=\sum_{y\in\Sigma}\Pr(Y=y\mid Z)\Pr(X_{i},Z_{i}\mid Y=y,Z)$ (By the law of total probability)
@@
Proof of Lemma D.2.
$\left|\Pr(Z_{i},Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Z)\Pr(Z_{j}\mid Z)\right|$
$=$ $\left|\Pr(Z_{i},Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Y,Z)\Pr(Z_{j}\mid Y,Z)+\Pr(Z_{i}\mid Y,Z)\Pr(Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Z)\Pr(Z_{j}\mid Z)\right|$
$\leq$ $\left|\Pr(Z_{i},Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Y,Z)\Pr(Z_{j}\mid Y,Z)\right|+\left|\Pr(Z_{i}\mid Y,Z)\Pr(Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Z)\Pr(Z_{j}\mid Z)\right|$ (Triangle inequality)
$\leq$ $\epsilon+\left|\Pr(Z_{i}\mid Y,Z)\Pr(Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Z)\Pr(Z_{j}\mid Z)\right|.$ (4.2)
@@
$\left|\Pr(Z_{i}\mid Y,Z)\Pr(Z_{j}\mid Y,Z)-\Pr(Z_{i}\mid Z)\Pr(Z_{j}\mid Z)\right|$
$=$ $\left|\Pr(Z_{i}\mid Y,Z)(\Pr(Z_{j}\mid Y,Z)-\Pr(Z_{j}\mid Z))+\Pr(Z_{j}\mid Z)(\Pr(Z_{i}\mid Y,Z)-\Pr(Z_{i}\mid Z))\right|$
$\leq$ $\left|\Pr(Z_{i}\mid Y,Z)(\Pr(Z_{j}\mid Y,Z)-\Pr(Z_{j}\mid Z))\right|+\left|\Pr(Z_{j}\mid Z)(\Pr(Z_{i}\mid Y,Z)-\Pr(Z_{i}\mid Z))\right|$ (Triangle inequality)
$\leq$ $\left|\Pr(Z_{j}\mid Y,Z)-\Pr(Z_{j}\mid Z)\right|+\left|\Pr(Z_{i}\mid Y,Z)-\Pr(Z_{i}\mid Z)\right|$
$=$ $\frac{1}{\Pr(Y\mid Z)}\;\left|\Pr(Z_{j},Y\mid Z)-\Pr(Y\mid Z)\;\Pr(Z_{j}\mid Z)\right|+\frac{1}{\Pr(Y\mid Z)}\left|\Pr(Z_{i},Y\mid Z)-\Pr(Y\mid Z)\;\Pr(Z_{i}\mid Z)\right|$
@@ D.2 Necessity of the Assumptions
We show that this score is strictly smaller than the expected score when agents report 0 if and only if $X_{i}=Z_{i}=0$ (and $X_{j}=Z_{j}=0$) and report $1$ otherwise. Let $\theta$ be the strategy such that $\theta(k,l|z)=0$ if and only if $k=l=0$ for any $z$. We can write down the joint distribution and the product of marginal distributions for $\hat{X}_{i}$ and $\hat{X}_{j}$.
$\Pr(\hat{X}_{i},\hat{X}_{j}\mid y,z)=\begin{pmatrix}2d&0\\ 0&1-2d\end{pmatrix}$
$\Pr(\hat{X}_{i}\mid y,z)\Pr(\hat{X}_{j}\mid y,z)=\begin{pmatrix}4d^{2}&2d(1-2d)\\ 2d(1-2d)&(1-2d)^{2}\end{pmatrix}\quad\text{for any $y,z\in\{0,1\}$.}$
@@
$=S_{i}(X_{i},X_{j}\mid Z)+8d^{2}.$
D.3 Information Monotonicity under Lazy-reporting Strategies
@@
A lazy-reporting strategy is a convex combination of truth-telling and a no-effort strategy. Therefore, it suffices to show that the expected score of truth-telling is strictly higher than that of any no-effort strategy.
A no-effort strategy is essentially a mixture of two types of strategy. First, $\mu_{i}(X_{i},Z_{i})$ is independent of both $X_{i}$ and $Z_{i}$. In this case, we have shown in the proof of Theorem 4.3 that the expected score is 0, strictly less than the expected score of truth-telling, which is $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$.
Second, $\mu_{i}(X_{i},Z_{i})$ depends only on $Z_{i}$; let $\mu_{i}(X_{i},Z_{i})=\hat{Z}_{i}$. Conditioned on $Z$, the Markov chain $Z_{i}\rightarrow\hat{Z}_{i}$ holds. Now, fix agent $j$’s lazy-reporting strategy and let her report be $\hat{X}_{j}$. By the data processing inequality shannon1948mathematical, we have:
@@
Next, suppose agent $i$ reporting $Z_{i}$, which is the score-maximizing lazy-reporting strategy regardless of agent $j$’s strategy. For any lazy-reporting strategy such that $\hat{X}_{j}=\nu_{j}(X_{j},Z_{j})$ that agent $j$ plays, applying the data processing inequality to agent $j$’s report yields:
@@
Therefore, 4.5 is sufficient for information monotonicity under lazy-reporting strategies. ∎
Appendix E Learning the Scoring Function
In Section 4.1, we assumed that the underlying scoring function $T_{\Delta}$ is known. For complex signal structures, especially when the signal space is large, we have to learn the scoring function from the data. This introduces a concern: if a significant fraction of agents rely on LLMs, the learned scoring function $T_{\tilde{\Delta}}$ might be manipulated in a way that benefits the LLM-generated reports. We mitigate this concern using the following proposition. Let $S_{i}^{T}(\theta_{i},\theta_{j})$ be the expected score computed under the scoring function $T$ when agents’ strategies are $\theta_{i}$ and $\theta_{j}$ respectively.
Proposition E.1.
Intuitively, the proposition holds because when $T_{\Delta}$ is known and agents are truth-telling, the expected score is $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$. However, if $T_{\Delta}$ is mislearned, the expected score is upper bounded by $I_{\text{TVD}}(\hat{X}_{i};\hat{X}_{j}\mid Z)$. Since our assumptions ensure that $Z_{i}$ and $Z_{j}$ are independent given $Z$, the data processing inequality guarantees that $I_{\text{TVD}}(\hat{X}_{i};\hat{X}_{j}\mid Z)<I_{\text{TVD}}(X_{i};X_{j}\mid Z)$.
E.1 implies that the expected score of the conditioned CA mechanism is maximized if every agent exerts full effort and reports $X_{i}$, while any strategy that manipulates the true scoring function will only (weakly) decrease the expected score. [^2] Here we assume that the principal can learn the correct $T_{\Delta}$ if everyone reports their full-effort signal.
Proof of E.1.